# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ 2.2.1 Equations solved
The symbols here have their usual meaning, and this set of equations is closed by the ideal gas equation of state $e=p/(\gamma-1)$. In the following sub-section, we show that cloud crushing simulations are degenerate in the choice of cloud density (hence the value of initial pressure).
2.2.2 De-dimensionalization of cloud-crushing in uniform wind
Figure 1: The initialization of a cloud crushing simulation with the choice of code units in Eqs. 2.2. This simulation setup has translational symmetry because the background wind is uniform. Therefore, the cloud can, in principle, be placed anywhere. However, we choose to place it a few cloud radii from the left edge of the simulation domain as the wind blows from left to right. This ensures that the initial bow shock and the cloud tail are reliably captured within the simulation domain.
In this section, we outline the degenerate parameters in the cloud-crushing problem. We can de-dimensionalize Eqs. 2.1 in units of cloud size $R_{\rm cl}$, cloud density $\rho_{\rm cl}$, and wind velocity $v_{\rm wind}$ (which can be conveniently expressed in terms of the wind Mach number $\mathcal{M}$). Therefore, the following are the choices of units for de-dimensionalizing Eqs. 2.1,
@@ 2.2.2 De-dimensionalization of cloud-crushing in uniform wind
where $T_{\rm cl}$ is cloud temperature, $p_{\rm wind}$ is the wind pressure, $\chi$ is the density contrast between the cloud and the wind (such that $\chi=\rho_{\rm cl}/\rho_{\rm wind}$), $\gamma$ is the adiabatic index, and $\mu$ is the mean particle mass (in units of proton mass). The constants have their usual meaning. Using Eqs. 2.2, we de-dimensionalize the hydrodynamic equations into the following (using tildes to represent de-dimensionalized quantitites),
$\frac{\partial\tilde{\rho}}{\partial\tilde{t}}+\tilde{\nabla}.\left(\tilde{\rho}\tilde{\textbf{v}}\right)$ $=0,$ (2.3a)
$\frac{\partial}{\partial\tilde{t}}\left(\tilde{\rho}\tilde{\textbf{v}}\right)+\tilde{\nabla}.\left(\tilde{\rho}\tilde{\textbf{v}}\otimes\tilde{\textbf{v}}+\tilde{p}\mathcal{I}\right)$ $=0,$ (2.3b)
$\frac{\partial}{\partial\tilde{t}}\left(\frac{1}{2}\tilde{\rho}\tilde{v}^{2}+\tilde{e}\right)+\tilde{\nabla}.\left[\left(\frac{1}{2}\tilde{\rho}\tilde{v}^{2}+\tilde{e}+\tilde{p}\right)\tilde{\textbf{v}}\right]$ $=-\zeta\tilde{n}_{H}^{2}\tilde{\Lambda}(\tilde{T}).$ (2.3c)
Here $\zeta=\frac{X_{H}\mu}{\gamma(\gamma-1)}\mathcal{M}^{-2}\chi^{-3/2}\left(\frac{t_{\rm cool,mix}}{t_{\rm cc}}\right)^{-1}\left[\frac{\Lambda\left(T_{0}\right)}{\Lambda\left(\sqrt{\chi}T_{\rm cl}\right)}\right]$ and $T_{0}=v_{0}^{2}\mu m_{p}/k_{B}=\mathcal{M}^{2}\gamma\chi T_{\rm cl}$, where $X_{H}$ is the total hyderogen mass fraction, and $\mu$ is particle mass (in units of proton mass). With $L_{0}$, $v_{0}$ and $\rho_{0}$ in Eqs. 2.2 as the choice of code units, the isobaric initial pressure $\tilde{p}$ in code units is
@@
The initialization of cloud-crushing simulations with the choice of code units in Eqs. 2.2 is illustrated in Fig. 1. The parameters that arise in the initialization, boundary condition, and the evolution equations are the only relevant parameters in the cloud-crushing problem. The rest of the parameters, on the other hand, that do not appear in the initial or boundary conditions, as well as in the evolution equations, are degenerate. The relevant parameters are, therefore, $\chi$, $\mathcal{M}$, $t_{\rm cool,mix}/t_{\rm cc}\propto R_{\rm thres}/R_{\rm cl}$, and $T_{\rm cl}$. [^2] The simulation parameters also depend implicitly on the cooling function $\Lambda(T)$ which goes in $\zeta$ defined earlier. As long as these parameters are fixed, the initial and boundary conditions, and the evolution (appropriately scaled) remain unchanged. For simulations without radiative cooling ($t_{\rm cool,mix}/t_{\rm cc}\rightarrow\infty$), the set of free parameters reduces to only the density contrast $\chi$ and the Mach number of the wind $\mathcal{M}$. It becomes degenerate for any arbitrary choice of both the initial pressure $p_{\rm wind}$ and the initial size of the cloud $R_{\rm cl}$ (see also Scannapieco & Brüggen 2015).
2.2.3 De-dimensionalization of cloud-crushing in CC85 wind
Figure 2: The initialization of a cloud crushing simulation with the choice of code units in Eqs. 2.5. Unlike ‘vanilla’ cloud crushing, here the seed cloud is embedded in a CC85 wind. This introduces a background structure to the wind and breaks the translational symmetry of the initial condition of the ‘vanilla’ cloud-crushing setup. In addition, the wind is directed radially outward instead of being plane-parallel. Refer to Section 2.2.3 for an explanation of each of the terms in this cartoon.
We extend this de-dimensionalization technique to simulations where clouds are seeded in expanding outflows, specifically the steadily expanding adiabatic outflows modeled by Chevalier & Clegg 1985 (CC85). The main difference between these setups and ‘vanilla’ cloud crushing is that the background wind conditions have a spatial gradient. Additionally, instead of being plane parallel, the wind is blown radially in our simulation. In this case, the most convenient choice of code units for de-dimensionalization is the wind condition at the initial position of the center of the cloud. In the de-dimensionalized form, the hydrodynamic equations remain the same as Eqs. 2.3. The choice of variables for de-dimensionalization is as follows
@@ 2.2.3 De-dimensionalization of cloud-crushing in CC85 wind
Here, $\mkern 1.5mu\overline{\mkern-1.5muv\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\rho\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ are dimensionless velocity and density solutions respectively of the CC85 profile at the initial position of the center of the cloud [^3] The CC85 density, pressure, and velocity profiles are de-dimensionalized using $\dot{M}^{3/2}\dot{E}^{-1/2}d_{\rm inj}^{-2}$, $\dot{M}^{1/2}\dot{E}^{1/2}d_{\rm inj}^{-2}$, and $\dot{M}^{-1/2}\dot{E}^{1/2}$ respectively. They are denoted by overbars. With this normalization, all the CC85 profiles exactly trace each other, with distance in units of $d_{\rm inj}$. $d_{\rm ini}$. The steady injection rates of energy and mass are $\dot{E}$ and $\dot{M}$ respectively within an injection region with radius $d_{\rm inj}$ (sometimes also called the sonic radius as the Mach number of the wind there is unity) in a CC85 wind. The distance from the center of the wind in the dimensionless CC85 profile $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu$ is expressed in terms of the injection radius $d_{\rm inj}$, such that $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}=d_{\rm ini}/d_{\rm inj}$. In cloud-crushing simulations, the Mach number of the wind at the initial position of the center of the cloud $\mathcal{M}_{\rm ini}$ is a free parameter. Therefore, inverting the CC85 profile using $\mathcal{M}_{\rm ini}$, we get $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, which is thus uniquely specified by $\mathcal{M}_{\rm ini}$ as the Mach number profile of the wind $\mathcal{M}(\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu)$ in CC85 is an invertible function. Further, utilizing the value of $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, we determine the dimensionless CC85 wind properties $\mkern 1.5mu\overline{\mkern-1.5mu\rho\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, $\mkern 1.5mu\overline{\mkern-1.5muv\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ and $\mkern 1.5mu\overline{\mkern-1.5mup\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ at the initial position of the cloud to set up the conversion between physical units and code units (see Eqs. 2.5). The distance of any arbitrary point in the simulation domain with respect to the injection radius (needed to set the CC85 wind at that point) is $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu=(d/d_{\rm inj})(d_{\rm ini}/d_{\rm inj})=(d/d_{\rm ini})\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, where $(d/d_{\rm ini})$ is the distance of the arbitrary point with respect to the initial position of the cloud in any unit (code/physical) and $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ is actually a function $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}(\mathcal{M}_{\rm ini})$ obtained from $\mathcal{M}_{\rm ini}$ as described earlier. We use this distance $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu$ to populate the background wind condition in the simulation domain using the CC85 profile. The initial condition in the code units is illustrated in Fig. 2.
[…]
@@ 2.2.4 Simulation parameters
The values of the simulation parameters are listed in Table 1. The wind properties are chosen close to the values used in Chevalier & Clegg 1985, appropriate for the outflow in the M82 galaxy. The cloud properties and the initial Mach number of the wind are chosen to be similar to the previous cloud-crushing simulations like Gronke & Oh 2018, 2020; Kanjilal et al. 2021. This choice of free (and degenerate) parameters enables meaningful comparisons with previous studies and helps isolate the effects introduced by incorporating the expanding background of the outflow in our simulations.

### v3

@@ 2.1.1 Equations solved
Here $t$ denotes time, $\mathcal{I}$ is the identity tensor and $\rho,\textbf{v},p,$ and $e$ denotes density, velocity, pressure and internal energy density of the gas respectively. $\Lambda(T)$ that appears in the sink term in the energy equation (Eq. 2.1c) is the radiative cooling rate and $n_{H}=\rho X_{H}/m_{H}$ is the total hydrogen number density, where $X_{H}$ is the hydrogen mass fraction in the universe and $m_{H}$ is mass of the hydrogen atom. This set of equations (Eqs. 2.1) is closed by the ideal gas equation of state $e=p/(\gamma-1)$, where $\gamma$ is the ratio of specific heats with constant pressure and that with constant volume.
2.1.2 De-dimensionalisation of cloud-crushing in uniform wind
Figure 1: The initialisation of a cloud-crushing simulation with the choice of code units in Eqs. $2.$ This simulation setup has translational symmetry because the background wind is uniform. Therefore, the cloud can, in principle, be placed anywhere. However, we choose to place it a few cloud radii from the left edge of the simulation domain as the wind blows from left to right. This ensures that the initial bow shock and the cloud tail are reliably captured within the simulation domain.
In this section, we outline the degenerate parameters in the cloud-crushing problem. We can de-dimensionalise Eqs. 2.1 in units of cloud size $R_{\rm cl}$, cloud density $\rho_{\rm cl}$, and wind velocity $v_{\rm wind}$ (which can be conveniently expressed in terms of the wind Mach number $\mathcal{M}$). Therefore, the following are the choices of units for de-dimensionalising Eqs. 2.1,
@@ 2.1.2 De-dimensionalisation of cloud-crushing in uniform wind
where $T_{\rm cl}$ is cloud temperature, $p_{\rm wind}$ is the wind pressure, $\chi$ is the density contrast between the cloud and the wind (such that $\chi=\rho_{\rm cl}/\rho_{\rm wind}$), $\gamma$ is the adiabatic index, and $\mu$ is the mean particle mass (in units of proton mass). The constants have their usual meaning. Using Eqs. 2.2, we de-dimensionalise the hydrodynamic equations into the following (using tildes to represent de-dimensionalised quantities),
[…]
Here $\zeta=\frac{X_{H}\mu}{\gamma(\gamma-1)}\mathcal{M}^{-2}\chi^{-3/2}\left(\frac{t_{\rm cool,mix}}{t_{\rm cc}}\right)^{-1}\left[\frac{\Lambda\left(T_{0}\right)}{\Lambda\left(\sqrt{\chi}T_{\rm cl}\right)}\right]$ and $T_{0}=v_{0}^{2}\mu m_{p}/k_{B}=\mathcal{M}^{2}\gamma\chi T_{\rm cl}$, where $X_{H}$ is the total hydrogen mass fraction, and $\mu$ is particle mass (in units of proton mass). With $L_{0}$, $v_{0}$ and $\rho_{0}$ in Eqs. 2.2 as the choice of code units, the isobaric initial pressure $\tilde{p}$ in code units is
@@
The initialisation of cloud-crushing simulations with the choice of code units in Eqs. 2.2 is illustrated in Fig. 1. The parameters that arise in the initialisation, boundary condition, and the evolution equations are the only relevant parameters in the cloud-crushing problem. The rest of the parameters, on the other hand, that do not appear in the initial or boundary conditions, as well as in the evolution equations, are degenerate. The relevant parameters are, therefore, $\chi$, $\mathcal{M}$, $t_{\rm cool,mix}/t_{\rm cc}\propto R_{\rm thres}/R_{\rm cl}$, and $T_{\rm cl}$. [^2] The simulation parameters also depend implicitly on the cooling function $\Lambda(T)$ which goes in $\zeta$ defined earlier. As long as these parameters are fixed, the initial and boundary conditions, and the evolution (appropriately scaled) remain unchanged. Therefore, cloud-crushing simulations are degenerate in the choice of cloud density (hence the value of initial pressure). For simulations without radiative cooling ($t_{\rm cool,mix}/t_{\rm cc}\rightarrow\infty$), the set of free parameters reduces to only the density contrast $\chi$ and the Mach number of the wind $\mathcal{M}$. It becomes degenerate for arbitrary choice of both the initial pressure $p_{\rm wind}$ and the initial size of the cloud $R_{\rm cl}$ (see also Scannapieco & Brüggen 2015).
2.1.3 De-dimensionalisation of cloud-crushing in CC85 wind
Figure 2: The initialisation of a cloud-crushing simulation with the choice of code units in Eqs. 2.5. Unlike ‘vanilla’ cloud-crushing, here the seed cloud is embedded in a CC85 wind. This introduces a background structure to the wind and breaks the translational symmetry of the initial condition of the ‘vanilla’ cloud-crushing setup. In addition, the wind is directed radially outward instead of being plane-parallel. Refer to Section 2.1.3 for an explanation of each of the terms in this cartoon.
We extend this de-dimensionalisation technique to simulations where clouds are seeded in expanding outflows, specifically the steadily expanding adiabatic outflows modelled by Chevalier & Clegg 1985 (CC85). The main difference between these setups and ‘vanilla’ cloud-crushing is that the background wind conditions have a spatial gradient. Additionally, instead of being plane parallel, the wind is blown radially in our simulation. In this case, the most convenient choice of code units for de-dimensionalisation is the wind condition at the initial position of the centre of the cloud. In the de-dimensionalised form, the hydrodynamic equations remain the same as Eqs. 2.3. The choice of variables for de-dimensionalisation is as follows
@@ 2.1.3 De-dimensionalisation of cloud-crushing in CC85 wind
Here, $\mkern 1.5mu\overline{\mkern-1.5muv\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\rho\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ are dimensionless velocity and density solutions respectively of the CC85 profile at the initial position of the centre of the cloud [^3] The CC85 density, pressure, and velocity profiles are de-dimensionalised using $\dot{M}^{3/2}\dot{E}^{-1/2}d_{\rm inj}^{-2}$, $\dot{M}^{1/2}\dot{E}^{1/2}d_{\rm inj}^{-2}$, and $\dot{M}^{-1/2}\dot{E}^{1/2}$ respectively. They are denoted by overbars. With this normalisation, all the CC85 profiles exactly trace each other, with distance in units of $d_{\rm inj}$. $d_{\rm ini}$. The steady injection rates of energy and mass are $\dot{E}$ and $\dot{M}$ respectively within an injection region with radius $d_{\rm inj}$ (sometimes also called the sonic radius as the Mach number of the wind there is unity in a CC85 wind). The distance from the centre of the wind in the dimensionless CC85 profile $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu=d/d_{\rm inj}$ is expressed in terms of the injection radius $d_{\rm inj}$, such that $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}=d_{\rm ini}/d_{\rm inj}$. In cloud-crushing simulations, the Mach number of the wind at the initial position of the centre of the cloud $\mathcal{M}_{\rm ini}$ is a free parameter. On inverting the CC85 profile using $\mathcal{M}_{\rm ini}$, we get the dimensionless initial position of the cloud $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$. It is uniquely specified by $\mathcal{M}_{\rm ini}$ as the Mach number profile of the wind $\mathcal{M}(\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu)$ in the CC85 solution is an invertible function, i.e., $d_{\rm ini}/d_{\rm inj}=\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}=\mathcal{M}^{-1}(\mathcal{M}_{\rm ini})$, where $\mathcal{M}^{-1}(\mathcal{M}_{\rm ini})$ is equivalently expressed as $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}(\mathcal{M}_{\rm ini})$ in Fig. 2. Further, utilising the value of $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ (obtained by inverting the Mach number profile), we determine the dimensionless CC85 wind properties $\mkern 1.5mu\overline{\mkern-1.5mu\rho\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, $\mkern 1.5mu\overline{\mkern-1.5muv\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ and $\mkern 1.5mu\overline{\mkern-1.5mup\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ at the initial position of the cloud required for setting up the conversion between physical units and code units (see Eqs. 2.5). The distance of an arbitrary point in the simulation domain with respect to the injection radius (needed to set the CC85 wind at that point) is $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu=(d/d_{\rm inj})(d_{\rm ini}/d_{\rm inj})=(d/d_{\rm ini})\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, where $(d/d_{\rm ini})$ is the distance of the arbitrary point with respect to the initial position of the cloud in any unit (code/physical) and $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$ is actually a function $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}(\mathcal{M}_{\rm ini})$ obtained from $\mathcal{M}_{\rm ini}$ as described earlier. We use this distance $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu$ to populate the background wind condition in the simulation domain using the CC85 profile. The initial condition in the code units is illustrated in Fig. 2.
The set of independent free parameters for the cloud-crushing simulations in a CC85 wind remains the same as those in ‘vanilla’ cloud-crushing, i.e., $\chi$, $\mathcal{M}_{\rm ini}$, $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ and $T_{\rm cl}$ except for an additional parameter $d_{\rm ini}/R_{\rm cl}$. This parameter $d_{\rm ini}/R_{\rm cl}$ comes from the initialisation of the CC85 wind using the dimensionless position of every fluid element in the simulation domain $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu=[(d/R_{\rm cl})/(d_{\rm ini}/R_{\rm cl})]\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}$, where $(d/R_{\rm cl})$ and $(d_{\rm ini}/R_{\rm cl})$ are positions in code units of any point in the simulation domain and the initial centre of the cloud, respectively. This wind profile, set by the free parameter $(d_{\rm ini}/R_{\rm cl})$, breaks the translation symmetry of ‘vanilla’ cloud-crushing simulations and introduces the spatially changing background. Naturally, more degenerate quantities arise here – $\dot{E}$, $\dot{M}$, $d_{\rm inj}$, and $R_{\rm cl}$, which are all absorbed in the conversion factors between the physical and code units (see Eqs. 2.5). Specific to our parameter sets (cf. Table 1), the implicit relation between the free parameters of cloud-crushing simulations and the corresponding degenerate parameters are discussed in the next sub-section.
2.2 Initial & boundary conditions
Our initial condition is a stationary, dense, spherical cloud of radius $R_{\rm cl}$ at a temperature of $T_{\rm cl}=4\times 10^{4}\ \rm K$ in the simulation domain. The cloud is initialised to be locally isobaric with the background wind. We fix the initial cloud position and wind properties (to either Set 1 or Set 2 as listed in Table 1) and change the cloud size by varying $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ to systematically explore the growth/destruction of clouds in an expanding background. Because we are fixing the wind properties and the initial position of the cloud (hence $\mathcal{M}_{\rm ini}$), the two independent parameters in these cloud-crushing simulations, namely $d_{\rm ini}/R_{\rm cl}$ and $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$, no longer remain independent (cf. Section 2.1.3). We adopt the mass and energy injection rates, as well as the injection radius, from the values listed in Table 1. Since these parameters are degenerate, multiple combinations are possible. To enable a meaningful comparison with previous cloud-crushing simulations, we select values such that the wind temperature at the initial position of the cloud is $\chi T_{\rm cl}$. Using this constraint, we invert the dimensionless CC85 profile to obtain $\mkern 1.5mu\overline{\mkern-1.5mud\mkern-1.5mu}\mkern 1.5mu_{\rm ini}=d_{\rm ini}/d_{\rm inj}$. Based on this setup, we then compute consistent values for the Mach number at the initial centre of the cloud $\mathcal{M}_{\rm ini}$ and the ratio $d_{\rm ini}/R_{\rm cl}$, where $R_{\rm cl}$ is determined by the chosen value of $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}$ (also listed in Table 1). However, it is worth noting that our results are broadly applicable and other combination of wind properties (degenerate parameters) can also produce the same set of $(\chi,\mathcal{M}_{\rm ini},T_{\rm cl},t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini},d_{\rm ini}/R_{\rm cl})$ which will have identical cloud-wind interaction.
Therefore, by construction, the density of the cloud is initialised to a contrast $\chi$ with the background wind at the position of the centre of the cloud. The background condition outside the cloud is determined by the steady spherically symmetric solution of a CC85 wind (Chevalier & Clegg 1985). Consequently, the cloud density has a slight variation from its centre to the edge approximately by a factor of $2(\gamma-1)/(d_{\rm ini}/R_{\rm cl})$, which corresponds to at most $5\%$ for the largest cloud in our parameter space (lower for smaller clouds). In addition, we have ensured that our computational domain always excludes the region within the injection radius (sonic point) of the CC85 wind. This simplifies our numerical scheme, as the fluid equations beyond the sonic point do not require the inclusion of any mass, momentum, or energy injection sources. We were able to maintain a stable and steady background wind profile in the entire simulation domain consistent with the CC85 solution by fixing the boundary condition to any particular CC85 wind profile.
@@ 2.2.1 Simulation parameters
The simulation parameters chosen in our study are listed in Table 1. The wind properties are chosen close to the values reported by Chevalier & Clegg 1985, appropriate for the outflow in the M82 galaxy. The cloud properties and the initial Mach number of the wind are chosen to be similar to the previous cloud-crushing simulations like Gronke & Oh 2018, 2020; Kanjilal et al. 2021. Although our results are broadly applicable, our choice of free (and degenerate) parameters enables meaningful comparisons with previous studies and helps isolate the effects introduced by incorporating the expanding background of the outflow in our simulations. Discussions related to convergence and resolution dependence of our simulations are presented in Appendix E.